A store purchased 70 toolboxes for a total cost of $912. It sold 45 of these toolboxes for $18 each, 10 for $15 and 15 for $12 each. How much gross profit was made on the entire transaction? What percent of cost was the markup? What percent of selling was the markup?

Answer by stanbon(75887)

A store purchased 70 toolboxes for a total cost of $912. It sold 45 of these toolboxes for $18 each, 10 for $15 and 15 for $12 each. How much gross profit was made on the entire transaction? What percent of cost was the markup? What percent of selling was the markup?
------------------
Revenue = 45*18 + 10*15 + 15*12 = 1140
---
Profit = 1140-912 = 228
------
Profit Percent markup = 228/912 = 25%
-----
Selling Percent markup = 1140/912 = 125%
------------
Cheers,
Stan H.